OpenStudy (anonymous):

the formula for the volume of a cone is v=1/3pir^2. find the radius to the nearest hundredth, of a cone with a hieght of 5 inches and a volume of 20inches^3

hero (hero):

1. Isolate r:\[V = \frac{\pi r^2h}{3}\]\[3V = \pi r^2 h\]\[\frac{3V}{\pi h} = r^2\]\[\sqrt{\frac{3V}{\pi h}} = r\] 2. Plug in given values for V and h: \[\sqrt{\frac{3(20)}{\pi (5)}} = r\] 3. Evaluate the left side using your calculator:

OpenStudy (anonymous):

what do you mean evaluate it ?

OpenStudy (anonymous):

is it 3.82in

OpenStudy (anonymous):

@Hero is that right

OpenStudy (anonymous):

or1.95in

hero (hero):

1.95 is correct

OpenStudy (anonymous):

ok thanks!!!!!!

OpenStudy (anonymous):

@Hero can u tell me if this is correct.